### v2

# Compressed code: the hidden effects of quantization and distillation on programming tokens

Abstract
Large Language Models (LLMs) have demonstrated exceptional code generation capabilities, yet their token-level mechanisms remain underexplored, particularly in compressed models. Through systematic analysis of programming language token representations, we characterize how programming languages are encoded in LLM tokenizers by analyzing their vocabulary distribution and keyword coverage patterns. We introduce a novel cold-start probability analysis method that provides insights into model behavior without requiring explicit prompts. Additionally, we present a comprehensive evaluation of how different model optimization techniques - including quantization, distillation, model scaling, and task-specific fine-tuning - affect token-level representations and code generation quality. Our experiments, supported by comprehensive probability distribution analysis and evaluation metrics, reveal critical insights into token-level behavior and provide empirically-validated guidelines for maintaining code generation quality under various optimization constraints. These findings advance both theoretical understanding of LLM code generation and practical implementation of optimized models in production environments.
The reproducible code for our work is available on GitHub: https://github.com/Egelvein/compressed_code.
1 Introduction
1.1 Motivation
Programming language understanding and code generation are critical applications of LLMs. Tokenization plays a crucial role in shaping model efficiency and performance. Byte Pair Encoding (Bostrom and Durrett, 2020) is widely used but often introduce inefficiencies, particularly in programming languages where syntax-specific tokens must be preserved. Poor tokenization can lead to increased sequence length, higher computational cost, and reduced model interpretability.
While LLMs excel at natural language tasks, their ability to handle programming languages remains dependent on how well their tokenizers preserve syntax and semantics. Many tokenizers are optimized for general text but fail to efficiently capture code structure, leading to suboptimal code completion and generation results (see Dagan and others 2024). Recent work has shown that tokenizer quality significantly impacts model performance across different languages and domains (Chelombitko et al. 2024), with specialized tokenizers demonstrating substantial improvements in compression ratios and representation efficiency (see Chelombitko and Komissarov 2024). An emerging alternative direction explores tokenizer-free architectures: the Meta-AI Byte Latent Transformer (see Pagnoni et al. 2024) demonstrates that byte-level models can match tokenization-based LLM performance at scale, suggesting that the relationship between tokenization and model capabilities may be more nuanced than previously assumed.
Industry adoption of LLM compression techniques such as quantization and distillation is rapidly increasing, with over 73% of enterprise deployments relying on compressed models. However, little research has been conducted on how compression impacts programming language tokenization. Our study aims to fill this gap by systematically analyzing the effects of compression on token representation and model performance.
Efficient model deployment requires balancing compression techniques with maintaining generation quality. Understanding how token probability distributions shift due to compression will provide insights into ensuring reliable code completion in real-world applications.
1.2 Background and Related Work
Recent research has focused on improving LLM performance for natural language tasks, yet few studies have delved into token-level behavior in code-specific contexts. Prior work on model compression has largely targeted natural language tasks, leaving a gap in our understanding of its effects on programming language tokens. Our work builds on and extends these findings by examining token vocabulary composition, generation probability distributions, and model confidence metrics specifically for programming languages.
Despite these advances, few studies delve deeply into the granularity of token representations for programming languages. Existing research often assesses model performance via downstream tasks using traditional benchmarks like HumanEval (Chen and others, 2021) or MBPP (Austin and others, 2021). However, the shift towards evaluating complex instruction following and repository-level reasoning exemplified by BigCodeBench (Zhuo et al., 2025) and SWE-bench (Jimenez et al., 2024) - highlights the need for deeper interpretability. These modern benchmarks reveal that high pass rates do not always correlate with structural understanding, yet few studies link these performance metrics to the granularity of token representations or probability evolution.
Common model compression strategies include quantization (reducing numerical precision), distillation (teacher-student paradigms), and pruning (removing weights or neurons). While compression methods are widely studied for natural language tasks, code-focused compression remains underexplored. Preliminary evidence suggests that quantization can degrade the ability to handle rare tokens (Ouyang and others, 2024), and distillation can lead to biased distributions if specialized tokens are not preserved.
By examining token probabilities pre- and post-compression, our work provides a clearer picture of why and where code completion might fail in constrained environments.
1.3 Our Contribution
Our primary contributions are threefold. First, we introduce a novel methodology for evaluating code-specific model capabilities through token probability analysis, including our ”cold start” metric that provides insights into model behavior without requiring extensive testing or benchmarking. This approach offers a new perspective on how models represent and process programming constructs, moving beyond traditional accuracy-based evaluations.
Second, we present surprising findings about the nature of code-specialized models, demonstrating that they do not develop fundamentally different token distributions compared to general language models. This insight challenges common assumptions about model specialization and suggests that the relationship between tokenization and code generation capabilities is more nuanced than previously understood.
Third, we provide the first systematic analysis of how compression techniques affect code generation capabilities at the token level. Our results reveal unexpected non-linear effects, where moderate quantization can actually improve the balance between programming keywords and special tokens, while distillation can significantly alter token probability distributions. These findings have immediate practical implications for deploying compressed models in production environments while maintaining code generation quality.
This research bridges the gap between theoretical understanding of language models and practical deployment considerations, offering both novel analytical tools and actionable insights for maintaining reliable code generation capabilities under resource constraints.
2 Methodology and Results
2.1 Choosing coding models and tokenizers
As a starting point in our study, we selected a set of top-performing open-source language models that demonstrated the best performance in coding tasks while also being publicly available in the Chatbot Arena benchmark (Chiang and others, 2024). The models included in our analysis are Qwen2.5-Coder, Qwen2.5 (Hui and others, 2024), Athene-V2-Chat (Nexusflow, 2024), DeepSeek-V2.5 (DeepSeek-AI et al., 2024a), DeepSeek-V3 (DeepSeek-AI et al., 2024b), DeepSeek-R1, and additionally DeepSeek-R1 distillations: DeepSeek-R1-Llama-70B, DeepSeek-R1-Llama-8B, DeepSeek-R1-Qwen-32B, DeepSeek-R1-Qwen-14B, DeepSeek-R1-Qwen-7B, DeepSeek-R1-Qwen-1.5B (DeepSeek-AI et al., 2025), and Llama 3.1 (Grattafiori and others, 2024).
Upon analyzing the tokenizer vocabularies of the selected models, we identified several instances where different models share identical tokenization schemes. Specifically, we found that Qwen2.5-Coder, Qwen2.5, Athene-V2-Chat, DeepSeek-R1-Qwen-32B, DeepSeek-R1-Qwen-14B, DeepSeek-R1-Qwen-7B, and DeepSeek-R1-Qwen-1.5B use the same vocabulary. Similarly, DeepSeek-R1 and DeepSeek-V3 share an identical vocabulary, indicating continuity in tokenization strategies across different versions of the DeepSeek series. Finally, we observed that Llama 3.1, DeepSeek-R1-Llama-70B, and DeepSeek-R1-Llama-8B use the same tokenizer vocabulary, which aligns with expectations given their architectural similarities. Thus, only three unique tokenizers remain for further analysis, which we will focus on in our study. From now on, we will refer to the tokenizers by the names of their respective models.
2.2 Keywords as language markers in tokenizers
Next, we considered the most popular programming languages on GitHub (Beuke, 2023), as this metric better reflects real-world usage and developer activity. Based on recent rankings, we selected the top ten programming languages: Python, Java, Go, JavaScript, C++, TypeScript, PHP, Ruby, C, and C#. Additionally, we included Rust and React, the last is a widely used JavaScript library for building user interfaces, as it has gained significant traction in code generation tasks using large language models. To assess how well these languages and frameworks are represented in the tokenizers, we adopted the following approach. For each programming language, we compiled a set of its reserved keywords, while for React, we collected key framework-specific keywords. These sets were then used as a basis for evaluating their presence in the tokenizers.
To further analyze how programming-related tokens are represented in three tokenizers, we compiled a dataset of reserved keywords from multiple programming languages (276 keywords) and examined their rankings.
Figure 1: Pipeline for analyzing programming language tokens in LLM tokenizers across three major models (DeepSeek-R1/V3, Qwen2.5, and Llama 3.1). The process analyzes vocabulary composition (128K-151K tokens), examines keyword representation across 12 programming languages (276 unique keywords), and categorizes tokens by frequency to assess language coverage.
2.3 Vocabulary Analysis
We aimed to determine how well language-specific tokens are represented in the tokenizers, following the process depicted in Figure 1. This task can be divided into two parts. The first step was to identify which tokens should be considered language-specific. To address this, we selected tokens that correspond to reserved keywords in each programming language. The second step involved evaluating the rank of each token within the tokenizer’s vocabulary. For every token associated with a programming language, we examined its ranking to assess its prominence in the tokenizer. Appendix Tables LABEL:keyword-token-rank-analysis and 2 provides an overview of the presence and relative ranking of these keywords across tokenizers. The ranking reflects how frequently a token appears within a tokenizer’s vocabulary, with lower ranks indicating higher frequency.
DeepSeek-R1 tokenizer consistently lags behind Qwen2.5-Coder and Llama 3.1 across all languages, likely because new tokens were added at the beginning of its vocabulary, shifting coding tokens to the right, or because its training dataset contained significantly less code than those used for Qwen and Llama 3.1. We also observe that modern, widely-used languages such as Python, TypeScript, Rust, Go, and C# are well-represented, with keyword coverage exceeding 90% in most cases. This aligns with their prevalence in contemporary software development and AI-assisted coding environments. In contrast, languages like C, C++, and React exhibit notably lower coverage, with React being the least represented. This suggests that these tokenizers may not be optimized for frameworks and lower-level system programming languages, potentially leading to inefficiencies in language model usage for these domains.
A key factor influencing token representation is the overlap between programming language keywords and natural language vocabulary. Languages like Python have many keywords that also exist in general vocabulary (e.g., return, in, with), making them more likely to be represented as single tokens due to their high frequency in pretraining corpora. In contrast, more domain-specific keywords, such as those in C, C++, and Rust, are rarely found in general text, leading to fragmentation in tokenization or complete absence from token vocabularies.
This observation suggests that languages with higher lexical overlap with natural language may be better understood by models, as their keywords are directly mapped to existing tokens. However, this can also introduce ambiguity, as these tokens were seen in different context during the model training. On the other hand, specialized programming keywords that do not intersect with natural language have the advantage of being unambiguous but may suffer from suboptimal tokenization, impacting model performance in tasks requiring precise syntactic understanding.
However, the mere presence or absence of a token in the tokenizer is not a sufficiently illustrative metric. A more informative measure is the ranking of the token within the tokenizer’s vocabulary. This ranking provides deeper insight into how different tokenizers prioritize language-specific tokens and how they distribute them within their vocabularies.
2.4 Token ranking metric
In addition to measuring keyword coverage, we analyzed the ranks assigned to programming language keywords in different tokenizers. Our findings reveal a systematic trend where DeepSeek-R1 assigns significantly higher ranks to keywords compared to Qwen2.5-Coder and Llama 3.1. Since rank in BPE-based tokenization is largely influenced by token frequency in the training corpus, this suggests that DeepSeek-R1 either has a different tokenization strategy or its training data contains fewer instances of these programming language keywords.
This discrepancy reinforces our earlier observation that DeepSeek-R1 is structurally different from the other tokenizers. A higher rank indicates that a keyword is tokenized less efficiently, meaning it may be split into multiple subword tokens rather than appearing as a single unit. This could impact model performance, particularly for programming tasks where tokenization granularity plays a crucial role in syntax comprehension. The results suggest that DeepSeek-R1 may not be as optimized for code-related tasks as Qwen2.5-Coder and Llama 3.1, which assign lower ranks to programming keywords and thus likely tokenize them more efficiently.
Additionally we checked the presence specific not word-releated tokens (Appendix Table 3). Llama 3.1 and Qwen2.5 have a significantly higher proportion of tokens containing programming-related symbols compared to DeepSeek-V3. This suggests that DeepSeek-V3 may have a more compressed or selective representation of syntax-related tokens, potentially impacting its efficiency in parsing and generating code.
One unexpected result was that we initially assumed coding-specific models would have distinct token distributions compared to general language models, with additional tokens specifically introduced during training to better represent programming languages. However, this turned out not to be the case. General language models and those designed for coding share the same set of tokens. This finding was surprising, as we expected programming-focused models to exhibit a tokenizer vocabulary more specialized for coding-related syntax and keywords.
Another unexpected finding is that a large portion of tokens in these tokenizers consists of special characters commonly used in programming and formatting rather than carrying direct semantic meaning. In Llama 3.1 - 14.6% of the vocabulary is composed of such tokens, in Qwen2.5 it is 12.1%, and in DeepSeek-V3 it is 5.1%. This suggests that a considerable fraction of the tokenizer is allocated to structural elements—such as brackets, indentation, and operators—rather than words or identifiers. This distribution could impact how efficiently models process programming languages and structured data, as these non-semantic tokens play a crucial role in syntax representation but do not directly contribute to meaning in the traditional linguistic sense.
2.5 A new metric to estimate coding abilities
The next step in our analysis stemmed from the observation that the difference in tokens between models was not as significant as we initially expected. This can be explained by the fact that, despite analyzing multiple models, they all rely on only three distinct tokenizers. To gain further insight, we introduced an additional metric to evaluate how programming-related tokens are represented in pre-trained models. Specifically, we examined the probability of generating these tokens when no prompt is provided—a scenario we refer to as the cold start. Under the cold start condition, we analyzed the individual probabilities of programming-related tokens appearing spontaneously, as well as their cumulative probability across different models. However, we were only able to compute this metric for the following models: DeepSeek-R1-Distill-Qwen-1.5B, DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-Qwen-14B, DeepSeek-R1-Distill-Qwen-32B, DeepSeek-R1-Distill-Llama-8B, DeepSeek-R1-Distill-Llama-70B, Qwen2.5-Coder-7B-Instruct from DeepSeek R1 repository.
2.5.1 Metrics Definition
To comprehensively evaluate the models’ coding capabilities, we introduced several key metrics:
  1. 1.
    Programming Keywords Probability (PKP): This metric represents the cumulative probability of generating programming language keywords (e.g., ’def’, ’import’, ’class’) in a cold start scenario. The PKP provides insight into the model’s baseline tendency to generate actual programming constructs.
  2. 2.
    Special Tokens Probability (STP): This metric measures the cumulative probability of generating special tokens commonly used in programming (e.g., parentheses, brackets, indentation markers). The STP helps understand the model’s structural coding awareness.
  3. 3.
    Keyword Average Probability (KAP): Calculated as the mean probability across all programming keywords, this metric allows us to assess the uniform distribution of keyword probabilities and identify potential biases toward specific programming constructs.
  4. 4.
    Natural Language Probability (NLP): This serves as a control metric, measuring the probability of generating common natural language tokens. It helps establish a baseline for comparison with programming-specific probabilities.
2.5.2 Results Analysis
Our analysis revealed several interesting patterns across the evaluated models (Appendix Table 4). First, we observed that the Special Tokens Probability consistently dominated the probability distribution, with values ranging from 0.1006 to 0.7246 across different models. This suggests a strong structural bias in the models’ architecture, favoring syntactic elements over semantic programming constructs. The Programming Keywords Probability showed more variation between models, with Qwen2.5-Coder-7B-Instruct demonstrating notably higher values (0.0902) compared to other models in the DeepSeek family (0.0011-0.0062). This significant difference suggests that architectural choices and training objectives may have a stronger influence on coding capabilities than model size alone. Interestingly, the Natural Language Probability remained relatively consistent across models (0.00013-0.0017), indicating that the models maintain a stable baseline for general language generation despite varying degrees of specialization for programming tasks. This observation provides a valuable reference point for understanding the models’ relative biases toward programming-specific tokens. When examining the Keyword Average Probability, we found that larger models did not necessarily exhibit higher probabilities for programming keywords. For instance, DeepSeek-R1-Distill-Qwen-1.5B showed a higher average probability (1.53E-05) compared to its larger counterpart DeepSeek-R1-Distill-Qwen-7B (2.78E-07), suggesting that model size may not be the primary factor in determining programming token representation.
2.5.3 Top Token Analysis
Further investigation of the most probable tokens in each category revealed distinct patterns across models. In programming keywords, commonly used terms like ’import’, ’def’, and ’void’ consistently appeared among the top-3 keywords, though their relative rankings varied between models. This suggests a common foundation in basic programming constructs while maintaining model-specific specializations. The analysis of top special tokens revealed particularly interesting patterns across the model families. The DeepSeek-R1-Distill-Qwen-1.5B showed a strong preference for basic syntax elements, with the newline character being its most probable special token. This suggests a fundamental understanding of code formatting structure. In contrast, the larger DeepSeek-R1-Distill-Qwen-7B favored XML-style tokens, potentially indicating a broader exposure to markup languages during training. Interestingly, model size appeared to influence the complexity of preferred special tokens. The DeepSeek-R1-Distill-Qwen-14B favored the C-style format specifier, while the 32B variant showed a preference for basic parentheses. The DeepSeek-R1-Distill-Llama-8B exhibited a unique pattern with the hash symbol as its top special token, suggesting possible optimization for Python-style comments or preprocessing directives. This variation in top special tokens across models provides insight into their architectural biases and potential specializations:
Syntax-focused models (e.g., DeepSeek-R1-Distill-Qwen-1.5B) with high probabilities for basic code formatting tokens Markup-oriented models (e.g., DeepSeek-R1-Distill-Qwen-7B) showing preference for XML/HTML-style tokens Language-specific models (e.g., DeepSeek-R1-Distill-Llama-8B) with special tokens commonly used in specific programming languages
The correlation between top programming keywords and special tokens further reinforces these specialization patterns. For instance, models with ’def’ and ’import’ among their top keywords often showed higher probabilities for Python-related special tokens, suggesting consistent language-specific training effects.
2.6 The Ultimate Pythonist’s Question: Tabs vs. Spaces
Beyond the traditional evaluation metrics, we explored a unique aspect of code generation: the models’ inherent preferences for different indentation styles. This analysis offers an empirical perspective on the famous ”tabs versus spaces” debate in Python programming. Our cold start probability analysis reveals several intriguing patterns (Appendix Table 5). First, the smallest model (DeepSeek-R1-Distill-Qwen-1.5B) shows an overwhelming preference for newlines (6.70E-02) over any indentation method, suggesting a focus on basic code structure rather than specific formatting. In contrast, larger models demonstrate more nuanced preferences, with the 32B variant showing relatively balanced probabilities between different indentation styles. Remarkably, across all models, spaces consistently outperform tabs in probability, with four-space indentation being particularly favored in larger models. The DeepSeek-R1-Distill-Llama-8B shows the most dramatic space-over-tab preference, with four-space indentation (8.71E-05) being approximately 60 times more probable than tabs (1.43E-06). This aligns with PEP 8 Python style guidelines, suggesting that these models have naturally absorbed common coding conventions during training. This finding adds an interesting dimension to our understanding of how models learn and represent programming conventions, showing that even without explicit programming, they develop clear preferences that mirror real-world coding practices.
2.7 Compression Impact
Next, we decided to focus on three key approaches commonly observed in the production deployment of models, beyond differences between architectures. Instead of comparing different models, we examined how a single family of models changes under three distinct processes:
  1. 1.
    Quantization: This process reduces the precision and dimensionality of numerical representations in model weights, effectively compressing the model while maintaining performance.
  2. 2.
    Model size: We examined models of different sizes (1.5B, 7B, 14B, and 32B) to assess how scale impacts performance and efficiency.
  3. 3.
    Model type: We considered three key variants—base models, instruct-tuned models, and code models - to evaluate their respective optimizations and intended use cases.
For our investigation, we focused exclusively on Qwen2.5 family of models, as they are open-source and provide well-structured series of models that vary systematically both in quantization levels and distillation stages. This allowed us to systematically analyze the effects of these transformations on programming language token representations (Appendix Table 6).
2.7.1 Quantization Effects
For a deeper understanding of how model compression affects coding abilities, we analyzed various quantization levels of the Qwen2.5-coder-7B model. It is interesting to note that the quantized Q8_0 model is very close to the original model. Our analysis revealed several critical patterns that challenge initial assumptions about model performance. A key finding emerged in the distribution between programming keywords and special tokens. While Q2_K showed a concerning ratio with high special token probability (0.4248) compared to keyword probability (0.0902), less aggressive quantization actually improved this balance. The Q4_K_S variant, for instance, demonstrated a more favorable distribution with reduced special token probability (0.2804) and increased keyword probability (0.1231). This observation aligns with recent findings on quantization robustness (Fang et al., 2025), which demonstrate that quantized models can exhibit superior resilience to noise perturbations compared to their full-precision counterparts. This shift is particularly significant because high special token probabilities indicate a model’s tendency to generate noise – repeated spaces, brackets, hashes, and other non-semantic elements that can lead to degenerate output patterns. For example, we observed tokens representing multiple repeated spaces, which could cause the model to get stuck in loops of meaningless character generation. The consistency in top programming keywords (’import’, ’package’, ’def’, ’from’) across quantization levels suggests that essential coding semantics are preserved. However, the reduced probability of special tokens in more aggressive quantization schemes might actually represent an improvement in generation quality, as it indicates a lower likelihood of producing syntactic noise.
2.8 Impact of Model Size and Instruction Tuning
Our analysis of the Qwen2.5 model family across different sizes (1.5B to 32B) and variants (base, instruct, and coder) revealed several critical patterns in code generation capabilities. The data presents a complex interplay between model size, specialization, and instruction tuning.
2.8.1 Base Model Scaling
In the base Qwen2.5 models, we observed a relatively stable pattern of keyword probabilities across different sizes (0.143 for 1.5B, 0.113 for 7B, 0.123 for 14B, and 0.142 for 32B). Notably, these models maintained a healthier ratio between programming keywords and special tokens, with special token probabilities remaining consistently lower (around 0.12-0.15). This suggests that base models develop a more balanced token distribution regardless of size.
2.8.2 Instruction Tuning Effects
A particularly interesting finding emerged in the impact of instruction tuning. Across all sizes, instruction-tuned variants showed a concerning trend toward increased special token probabilities. This effect was most visible in the 7B-Coder-Instruct model, where special token probability reached 0.449 – potentially indicating an increased risk of generating syntactic noise. The most striking degradation appeared in smaller instruction-tuned models. For instance, the Qwen2.5-Coder-1.5B-Instruct showed a drop in keyword probability (0.0003) compared to its base version (0.121), while maintaining a relatively high special token probability. This suggests that instruction tuning might be particularly challenging for smaller models to maintain code generation quality. The most significant difference we observe, aside from the fact that the Coder 1.5-Instruct model differs the most, is that it has completely different keywords. It’s unclear what happened to it, but it stands out considerably. In other words, there is one model where entirely different keywords suddenly appear at the top of the output. Additionally, its Keywords Probability is significantly lower.
2.8.3 Size-Dependent Patterns
The base 32B model maintained strong keyword probabilities (0.142) with balanced special token distribution The Coder-32B version showed the highest keyword probability (0.173) among all variants However, its instruction-tuned version (Coder-32B-Instruct) exhibited a significant degradation (0.027), suggesting that even large models aren’t immune to instruction tuning artifacts. These findings suggest that while larger models generally handle code generation better, instruction tuning can introduce unexpected biases toward special token generation, potentially compromising code generation quality regardless of model size.
2.9 Difference with distilled models
The comparison between original Qwen models and their distilled DeepSeek-R1 counterparts reveals striking differences in token probability distributions, suggesting significant transformations in model behavior through the distillation process. Most notably, the distilled models show lower programming keyword probabilities compared to their original counterparts. For instance, while Qwen2.5-1.5B maintains a healthy keyword probability of 0.143, its distilled version (DeepSeek-R1-Distill-Qwen-1.5B) drops to merely 0.0042 – a reduction of nearly 97%. This pattern persists across all model sizes, with the 32B variant showing a similar decrease from 0.142 to 0.0019. Even more concerning is the shift in special token probabilities. The distilled 1.5B model shows an increase in special token probability (0.7246) compared to its original version (0.125), suggesting a strong bias toward generating syntactic noise. This indicates that the distillation process might be inadvertently amplifying the model’s tendency to produce non-semantic tokens while diminishing its ability to generate meaningful programming constructs. The transformation is also evident in the models’ preferred tokens. While original Qwen-family models consistently favor semantic programming keywords (’import’, ’package’, ’from’), the distilled versions show a shift toward more basic constructs (’void’, ’int’, ’is’) and exhibit higher probabilities for formatting-related special tokens. This suggests that the distillation process, while potentially optimizing for model size and speed, might be compromising the models’ ability to generate coherent code. This degradation mirrors the theoretical findings (Dohmatob et al., 2024) regarding ’model collapse’. Their analysis demonstrates that models trained on synthetic outputs (as in the distillation process) tend to lose variance in their probability distributions, effectively pruning the vocabulary tails and converging onto a narrow set of high-frequency tokens. It raises important considerations about the trade-offs involved in model distillation, particularly for code generation tasks where maintaining a balance between semantic and syntactic elements is crucial.
2.10 Temperature effect
An interesting aspect of text generation is that the raw logits produced by a model can differ significantly from what the end user actually sees. This discrepancy arises due to post-processing techniques and sampling parameters, particularly temperature scaling, which can be implemented quite differently across models.
Temperature controls the level of randomness in token selection—lower values make the model more deterministic, while higher values increase diversity in output. However, different models may implement temperature scaling in non-uniform ways, leading to variations in how tokens are sampled even when given the same logits.
3 Conclusions
In this work, we present the first systematic token-level analysis of Large Language Models’ behavior in code generation tasks, with particular emphasis on the effects of compression techniques such as quantization and distillation on programming construct representations. Our results make several significant contributions to the field of code-generating language models. We introduce a novel analytical framework that includes a ”cold start” probability metric and a comprehensive set of quantitative measures for evaluating model behavior. These metrics - Programming Keywords Probability, Special Tokens Probability, Keyword Average Probability, and Natural Language Probability - provide new insights into models’ code generation capabilities without requiring extensive testing.
Through systematic application of these metrics, we demonstrate that token probability distributions at initialization can serve as reliable predictors of a model’s coding abilities. Our detailed characterization of tokenizer behavior across popular programming languages reveals several unexpected findings. Most notably, we discover that code-specialized models do not develop distinct token distributions compared to general language models, challenging common assumptions about model specialization. Furthermore, we identify significant differences in token-level behavior between models even when they share common vocabularies, suggesting that architectural choices and training objectives may have a stronger influence on coding capabilities than previously understood.
Through extensive experimentation with compression techniques, we uncover non-linear effects in model behavior. Particularly noteworthy is our finding that moderate quantization can actually improve the balance between programming keywords and special tokens, while more aggressive compression through distillation significantly alters token probability distributions, reducing programming keyword probabilities by up to 97% while increasing special token probabilities. We also identify concerning patterns in instruction-tuned models, particularly in smaller variants, where tuning can lead to decreased programming keyword probabilities. These findings not only advance our theoretical understanding of token-level mechanisms in code-generating language models but also provide practical guidelines for deploying compressed models in production environments. Our work establishes clear metrics and methodologies for evaluating the trade-offs between model compression and code generation capabilities, enabling more informed decisions in resource-constrained environments while maintaining reliable code generation quality.
Appendix A Appendix
Table 1: Keyword Token Rank Analysis. Columns include number of languages using each keyword, minimum rank, median rank, and ranks across tokenizers (Qwen2.5, DeepSeek-V3, Llama 3.1).
Keyword # Langs Min Rank Median Rank # Tokenizers Qwen2.5 DeepSeek-V3 Llama
in 6 258.0 258.0 3 258 261 258
or 4 269.0 269.0 3 269 272 269
is 2 278.0 285.0 3 285 278 285
as 5 300.0 300.0 3 300 306 300
if 11 333.0 333.0 3 333 394 333
int 4 396.0 396.0 3 396 650 396
end 1 408.0 408.0 3 408 523 408
out 1 411.0 412.0 3 411 606 412
and 4 437.0 438.0 3 437 458 438
import 5 474.0 475.0 3 474 1897 475
this 5 574.0 576.0 3 574 3779 576
return 11 689.0 693.0 3 689 3916 693
self 2 721.0 726.0 3 721 2161 726
def 2 750.0 755.0 3 750 3465 755
key 1 792.0 798.0 3 792 4989 798
use 2 810.0 817.0 3 810 3103 817
public 6 888.0 898.0 3 888 3978 898
string 1 917.0 928.0 3 917 4463 928
new 6 931.0 943.0 3 931 2839 943
var 4 947.0 959.0 3 947 5241 959
using 2 970.0 985.0 3 970 4079 985
include 1 997.0 1012.0 3 997 5211 1012
void 6 1004.0 1019.0 3 1004 6483 1019
lock 1 1023.0 1039.0 3 1023 10566 1039
const 8 1024.0 1040.0 3 1024 3949 1040
class 8 1040.0 1058.0 3 1040 3767 1058
ref 3 1097.0 1116.0 3 1097 5044 1116
let 3 1149.0 1169.0 3 1149 1775 1169
struct 5 1235.0 1257.0 3 1235 8325 1257
type 2 1313.0 1337.0 3 1313 4611 1337
print 1 1350.0 1374.0 3 1350 3098 1374
from 1 1499.0 1527.0 3 1499 5356 1527
else 11 1503.0 1531.0 3 1503 9267 1531
export 3 1533.0 1562.0 3 1533 7449 1562
try 8 1539.0 1568.0 3 1539 26976 1568
function 3 1688.0 1723.0 3 1688 8701 1723
object 1 1700.0 1735.0 3 1700 10325 1735
package 4 1722.0 1757.0 3 1722 7249 1757
select 1 1742.0 1779.0 3 1742 13933 1779
char 4 1762.0 1799.0 3 1762 7526 1799
true 7 1866.0 1904.0 3 1866 11476 1904
not 3 1921.0 1962.0 3 1921 2869 1962
for 11 1958.0 2000.0 3 1958 2251 2000
static 8 1978.0 2020.0 3 1978 15514 2020
private 6 1996.0 2039.0 3 1996 12339 2039
box 1 2011.0 2054.0 3 2011 7353 2054
context 1 2147.0 2196.0 3 2147 15689 2196
map 1 2186.0 2235.0 3 2186 10865 2235
Component 1 2189.0 2238.0 3 2189 10220 2238
assert 2 2207.0 2256.0 3 2207 12609 2256
signed 2 2215.0 2239.5 2 2215 - 2264
namespace 3 2231.0 2280.0 3 2231 10972 2280
default 8 2258.0 2309.0 3 2258 14979 2309
endif 1 2330.0 2384.0 3 2330 26297 2384
state 1 2454.0 2513.0 3 2454 9395 2513
uint 1 2496.0 2557.0 3 2496 34671 2557
True 1 2514.0 2575.0 3 2514 10634 2575
mod 1 2593.0 2658.0 3 2593 5158 2658
bool 3 2641.0 2707.0 3 2641 20600 2707
go 1 2710.0 3346.0 3 3346 2710 3427
func 1 2830.0 2900.0 3 2830 12642 2900
where 1 2870.0 2940.0 3 2870 4779 2940
NaN 4 2921.0 2994.0 3 2921 16042 2994
do 9 2982.0 3055.0 3 2982 4016 3055
echo 1 3047.0 3123.0 3 3047 18843 3123
event 1 3087.0 3163.0 3 3087 24875 3163
base 1 3152.0 3231.0 3 3152 14707 3231
then 1 3391.0 3473.0 3 3391 9594 3473
move 1 3397.0 3479.0 3 3397 47632 3479
while 10 3472.0 3556.0 3 3472 8848 3556
params 1 3519.0 3603.0 3 3519 20579 3603
next 1 3600.0 3684.0 3 3600 6695 3684
unsigned 2 3626.0 3710.0 3 3626 35536 3710
float 4 3649.0 3733.0 3 3649 15891 3733
byte 2 3782.0 3867.0 3 3782 23535 3867
false 7 3849.0 3934.0 3 3849 19836 3934
auto 2 3902.0 3989.0 3 3902 18414 3989
chan 1 3991.0 5658.0 3 5658 3991 5776
False 1 4049.0 4139.0 3 4049 15444 4139
None 1 4064.0 4155.0 3 4064 19906 4155
with 3 4197.0 4291.0 3 4197 6135 4291
template 1 4214.0 4308.0 3 4214 27793 4308
extern 4 4301.0 4399.0 3 4301 76215 4399
require 1 4310.0 4408.0 3 4310 22716 4408
double 4 4331.0 4429.0 3 4331 20563 4429
module 1 4352.0 4450.0 3 4352 20038 4450
delete 3 4542.0 4644.0 3 4542 23565 4644
END 1 4689.0 4794.0 3 4689 16860 4794
props 1 4761.0 4866.0 3 4761 27032 4866
long 4 4825.0 4930.0 3 4825 9938 4930
interface 6 4970.0 5077.0 3 4970 23964 5077
inline 2 5057.0 5167.0 3 5057 46900 5167
typedef 2 5286.0 5399.0 3 5286 29218 5399
case 9 5638.0 5756.0 3 5638 11675 5756
protected 6 5764.0 5883.0 3 5764 40252 5883
children 1 5864.0 5988.0 3 5864 20434 5988
boolean 1 6117.0 6245.0 3 6117 36541 6245
register 2 6343.0 6477.0 3 6343 28766 6477
match 2 6347.0 6481.0 3 6347 50374 6481
impl 1 6383.0 6517.0 3 6383 31281 6517
pass 1 6385.0 6519.0 3 6385 9762 6519
mut 1 6984.0 7129.0 3 6984 30293 7129
throw 6 7119.0 7265.0 3 7119 42117 7265
catch 6 7173.0 7320.0 3 7173 41163 7320
begin 1 7265.0 7413.0 3 7265 8277 7413
render 1 7322.0 7472.0 3 7322 28324 7472
checked 1 7549.0 7702.0 3 7549 47569 7702
async 2 7692.0 7847.0 3 7692 41072 7847
operator 2 7884.0 8043.0 3 7884 20178 8043
throws 1 8100.0 8262.0 3 8100 61101 8262
nil 1 8385.0 8551.0 3 8385 68782 8551
short 4 8676.0 8846.0 3 8676 31150 8846
foreach 2 8808.0 8984.0 3 8808 41810 8984
fn 2 8822.0 8998.0 3 8822 28952 8998
break 11 8960.0 9137.0 3 8960 15485 9137
enum 7 9018.0 9195.0 3 9018 47002 9195
typename 1 9031.0 9208.0 3 9031 50578 9208
override 2 9199.0 9380.0 3 9199 44462 9380
when 1 9309.0 9493.0 3 9309 21123 9493
super 5 9522.0 9712.0 3 9522 27196 9712
continue 10 9534.0 9726.0 3 9534 45911 9726
pub 1 9585.0 9780.0 3 9585 38744 9780
del 1 9588.0 9783.0 3 9588 18421 9783
range 1 9669.0 9866.0 3 9669 16801 9866
global 2 9752.0 9951.0 3 9752 31754 9951
typeof 5 10222.0 10433.0 3 10222 73890 10433
sizeof 3 10318.0 10531.0 3 10318 58850 10531
internal 1 10481.0 10701.0 3 10481 52430 10701
loop 1 10498.0 10719.0 3 10498 42938 10719
asm 1 10530.0 10753.0 3 10530 18969 10753
friend 1 10701.0 10931.0 3 10701 26736 10931
await 3 11421.0 11675.0 3 11421 62743 11675
except 1 11683.0 11945.0 3 11683 32812 11945
final 4 11822.0 12085.0 3 11822 41551 12085
priv 1 11887.0 12151.0 3 11887 78737 12151
Self 1 12092.0 12363.0 3 12092 37400 12363
elif 1 12458.0 12740.0 3 12458 53358 12740
lambda 1 12935.0 13077.0 3 12935 13077 13231
React 1 14799.0 15143.0 3 14799 54635 15143
alias 1 14956.0 15305.0 3 14956 88739 15305
extends 4 15231.0 15588.0 3 15231 83878 15588
undef 1 16135.0 16331.0 2 16135 - 16527
union 2 16192.0 16588.0 3 16192 32099 16588
abstract 4 16249.0 16647.0 3 16249 42149 16647
switch 8 17338.0 17790.0 3 17338 46941 17790
declare 1 18471.0 18978.0 3 18471 73780 18978
raise 1 18704.0 19223.0 3 18704 55003 19223
provider 1 19979.0 20576.0 3 19979 96058 20576
clone 1 19982.0 20579.0 3 19982 102131 20579
hook 1 20873.0 21543.0 3 20873 80513 21543
fixed 1 22021.0 22795.0 3 22021 77769 22795
endforeach 1 22095.0 22871.0 3 22095 122655 22871
readonly 2 22569.0 22977.0 2 22569 - 23385
decimal 1 23289.0 24170.0 3 23289 103814 24170
virtual 3 25668.0 26752.0 3 25668 76356 26752
elseif 1 26016.0 26561.5 2 26016 - 27107
ensure 1 27289.0 28389.0 3 27289 94747 28389
volatile 4 27307.0 27857.0 2 27307 - 28407
restrict 1 27898.0 28448.0 2 27898 - 28998
delegate 1 28227.0 28777.0 2 28227 - 29327
goto 5 28535.0 29635.0 3 28535 105723 29635
_Generic 1 29085.0 29635.0 2 29085 - 30185
trait 2 29432.0 29982.0 2 29432 - 30532
yield 6 29696.0 30246.0 2 29696 - 30796
native 1 29738.0 30838.0 3 29738 109720 30838
mutable 1 30473.0 31023.0 2 30473 - 31573
implicit 1 30940.0 31490.0 2 30940 - 32040
unchecked 1 31684.0 32234.0 2 31684 - 32784
macro 1 32606.0 33706.0 3 32606 116634 33706
ulong 1 32832.0 33382.0 2 32832 - 33932
BEGIN 1 37588.0 38688.0 3 37588 104472 38688
portal 1 37953.0 39053.0 3 37953 104008 39053
unsafe 2 38157.0 38707.0 2 38157 - 39257
until 1 38730.0 39830.0 3 38730 56657 39830
unless 1 38770.0 39870.0 3 38770 69190 39870
finally 6 39176.0 40276.0 3 39176 121229 40276
ushort 1 40375.0 40925.0 2 40375 - 41475
nullptr 2 41132.0 41682.0 2 41132 - 42232
requires 1 41375.0 42475.0 3 41375 101936 42475
fragment 1 42202.0 42752.0 2 42202 - 43302
constexpr 2 42281.0 42831.0 2 42281 - 43381
dyn 1 43085.0 43635.0 2 43085 - 44185
retry 1 44848.0 45398.0 2 44848 - 45948
jsx 1 45290.0 45840.0 2 45290 - 46390
consumer 1 46764.0 47864.0 3 46764 114495 47864
lazy 1 49013.0 49563.0 2 49013 - 50113
useState 1 55670.0 56220.0 2 55670 - 56770
crate 1 61711.0 62261.0 2 61711 - 62811
defer 1 62095.0 62645.0 2 62095 - 63195
redo 1 63561.0 64661.0 3 63561 127260 64661
elsif 1 65967.0 66517.0 2 65967 - 67067
concept 1 68487.0 69037.0 2 68487 - 69587
xor 2 71311.0 72411.0 3 71311 114718 72411
decltype 1 74364.0 74914.0 2 74364 - 75464
sealed 2 75940.0 76490.0 2 75940 - 77040
_Bool 1 79948.0 80498.0 2 79948 - 81048
typeid 1 88342.0 88892.0 2 88342 - 89442
compl 1 92843.0 92843.0 1 - 92843 -
explicit 2 93632.0 94182.0 2 93632 - 94732
implements 4 93958.0 94508.0 2 93958 - 95058
callable 1 95192.0 95742.0 2 95192 - 96292
Table 1: Keyword Token Rank Analysis. Columns include number of languages using each keyword, minimum rank, median rank, and ranks across tokenizers (Qwen2.5, DeepSeek-V3, Llama).
Table 2: Coverage of Programming Language Keywords in Tokenizer Vocabularies. A comparison of how well different tokenizers represent reserved keywords from various programming languages. Percentages indicate the proportion of keywords found in each tokenizer’s vocabulary.
Language Keywords Qwen2.5-Coder (%) DeepSeek-R1 (%) Llama (%)
C 59 71.2 59.3 71.2
C# 77 97.4 84.4 97.4
TypeScript 46 95.7 91.3 95.7
Ruby 41 87.8 78.0 87.8
PHP 62 85.5 75.8 85.5
Rust 51 96.1 86.3 96.1
JavaScript 46 95.7 91.3 95.7
Java 51 90.2 84.3 90.2
Python 35 97.1 94.3 97.1
Go 25 96.0 92.0 96.0
React 30 56.7 43.3 56.7
C++ 93 73.1 64.5 73.1
A full list of absent keywords in tokens: _Complex, static_assert, thread_local, _Thread_local, _Imaginary, _Alignas, _Alignof, alignas, _Decimal64, _Decimal128, _BitInt, _Atomic, _Static_assert, typeof_unqual, alignof, _Noreturn, _Decimal32, sbyte, stackalloc, debugger, instanceof, LINE, FILE, ENCODING, defined?, rescue, enddeclare, endfor, endswitch, endwhile, include_once, insteadof, require_once, yield from, become, unsized, permits, transient, strictfp, synchronized, nonlocal, fallthrough, reconciliation, useLayoutEffect, strictMode, useEffect, useImperativeHandle, errorBoundary, useCallback, suspense, useContext, useRef, useMemo, useReducer, virtualDOM, const_cast, bitor, constinit, not_eq, and_eq, noexcept, char8_t, xor_eq, co_yield, consteval, char16_t, bitand, or_eq, co_await, reinterpret_cast, wchar_t, static_cast, co_return, dynamic_cast, char32_t
Table 3: Proportion of Tokens Containing Programming-Specific Special Characters in Tokenizer Vocabularies. The number and percentage of tokens that contain programming-related symbols across three major tokenizers.
Tokenizer Tokens with Special Chars Total Tokens Percentage (%)
Llama 18,719 128,000 14.6
DeepSeek-V3 6,585 128,000 5.1
Qwen2.5 18,454 151,643 12.1
Table 4: Comparison of Tokenizer Statistics Across Models. This table presents keyword probability, special token probability, and other relevant statistics for various tokenizers.
Model KeyW Prob Spec tok Prob KeyW Avg Prob Spec tok Avg Prob NL prob Top-3 KeyW Top-3 Spec
[…]
R1-Qwen-14B 0.0011 0.1082 3.97E-06 5.36E-06 0.00013 is, in, use %c
[…]
R1-Llama-8B 0.0062 0.1006 2.25E-05 4.98E-06 0.0004 def, import, from #
Table 5: Comparison of Formatting Token Probabilities Across Models. This table presents tabulation, new line, and space token probabilities for various tokenizers.
Model Tab New line Two spaces Four spaces
R1-Qwen-1.5B 3.58E-07 6.70E-02 4.47E-06 1.01E-06
R1-Qwen-7B 4.29E-06 5.62E-05 5.90E-06 4.53E-06
R1-Qwen-14B 2.09E-06 6.59E-05 9.18E-06 1.48E-05
R1-Qwen-32B 2.03E-06 0.04132080078 0.0002186 7.99E-05
R1-Llama-8B 1.43E-06 9.95E-05 1.75E-05 8.71E-05
Table 6: Comparison of Tokenizer Statistics Across Models. This table presents keyword probability, special token probability, and other relevant statistics for various tokenizers.
Model KeyW Prob Spec tok Prob KeyW Avg Prob Spec tok Avg Prob NL prob Top-3 KeyW Top-3 Spec
Q2_K 0.0902 0.4248 0.0003 2.10E-05 0.0017 import, package, def **, #, //
Q3_K_L 0.1115 0.3289 0.0004 1.63E-05 0.0025 import, package, from **, #, //
Q3_K_M 0.1198 0.3320 0.0004 1.64E-05 0.0026 import, package, def **, #, //
Q3_K_S 0.0947 0.4147 0.0003 2.05E-05 0.0024 import, package, def **, #, //
Q4_0 0.0873 0.3353 0.0003 1.66E-05 0.0015 import, package, const **, #, //
Q4_1 0.0985 0.2821 0.0004 1.40E-05 0.0020 import, package, def **, #, //
Q4_K_M 0.1068 0.2900 0.0004 1.44E-05 0.0021 import, package, from **, #, //
Q4_K_S 0.1231 0.2804 0.0004 1.39E-05 0.0024 import, package, def **, #, //
Q5_0 0.1081 0.2830 0.0004 1.40E-05 0.0022 import, package, from **, #, //
Q5_1 0.1108 0.2867 0.0004 1.42E-05 0.0020 import, package, from **, #, //
Q5_K_M 0.1093 0.2929 0.0004 1.45E-05 0.0019 import, package, from **, #, //
Q5_K_S 0.1120 0.2928 0.0004 1.45E-05 0.0019 import, package, from **, #, //
Q6_K 0.1083 0.3011 0.0004 1.49E-05 0.0020 import, package, from **, #, //
Q8_0 0.1070 0.3025 0.0004 1.50E-05 0.0019 import, package, from **, #, //
Qwen2.5-1.5B 0.1434 0.1254 0.0005 6.22E-06 0.0043 import, package, from #, #include, <<
[…]
Qwen2.5-Coder-1.5B 0.1215 0.3104 0.0004 1.54E-05 0.0005 import, package, public **, #, /*
Qwen2.5-Coder-1.5B-Inst 0.0003 0.0262 1.10E-06 1.30E-06 5.47E-05 or, if, int ., ,, /
Qwen2.5-7B 0.1131 0.1470 0.0004 7.28E-06 0.0049 import, public, from #, #!/, #include
Qwen2.5-7B-Inst 0.1399 0.1644 0.0005 8.15E-06 0.0042 import, public, def #, <<, //
Qwen2.5-Coder-7B 0.1080 0.2985 0.0004 1.48E-05 0.0019 import, package, from **, #, //
Qwen2.5-Coder-7B-Inst 0.2572 0.4495 0.0009 2.23E-05 0.0007 import, package, const **, #, //
Qwen2.5-14B 0.1232 0.1249 0.0004 6.19E-06 0.0068 import, from, package #, #!/, <<
Qwen2.5-14B-Inst 0.0758 0.1062 0.0003 5.26E-06 0.0024 import, package, public #, <<, #!/
Qwen2.5-Coder-14B 0.0857 0.2587 0.0003 1.28E-05 0.0019 import, package, from **, #, //
Qwen2.5-Coder-14B-Inst 0.0108 0.2361 3.91E-05 1.17E-05 0.0026 package, import, public /, -, :